A Bibliography of Publications of Alessandro Saffiotti

Alessandro Saffiotti IRIDIA, Universite Libre de Bruxelles

50 av. F. Roosevelt, CP 194/6

Brussels Belgium 1050

Tel: +32 2 650 2729 FAX: +32 2 650 2715

E-mail: asaffio@ulb.ac.be (Internet)

13 October 2017 Version 1.09

Abstract

This bibliography records publications of Alessandro Saffiotti.

Title word cross-reference

4th [VY93].

’92 [VY93]. ’93 [Ano93a, CKM93, Ral94]. ’95 [DvLV96, IEE95]. ’96 [JPD96].

Acquaintance [SS88a, SS89]. advanced [VY93]. Advances [YKF94]. Agents [Saf91a]. AI [Saf87, Saf88]. ambiguity [KBB^{+}08]. Amsterdam [DvLV96]. anchor [KBB^{+}08]. Anchoring [CS99, KBB^{+}08]. Antibodies [CCM^{+}85]. application [CKS10]. applications [JPD96, YLZ95]. Approach [JS02, KBB^{+}08, SKR95, SS97]. approaches [CKM93]. approximate [Saf96]. April [IEE93]. Architecture [KMRS97]. Artificial [Ano93b, BHKL91, LS85, BH95, PS93, Ral94, VY93]. Assistant [CCM^{+}85]. August [BH95, Ral94]. Automatic [SU92]. Autonomous [RKS93, Saf93, Saf97b, Saf97c, Saf98a, Saf99, RSK95, SRK93c, SW95a, SW95b, Saf97a, Saf98c, Saf98b]. Autonomy [KMRS97]. aware [CKS10]. awareness [SS88a, SS89].

Based [VY93, BDL^{+}98, SW95a, SW96]. behavior [Saf97a]. Belief [BSS95a, BSS95b, BSS00, Saf91a, Saf92, SHS^{+}91, BS99, SS88a]. Believers [SS88b]. Best [PPSW99]. better [Saf88]. Blending [SRK93a]. Burlingame [Ano93a].

CA [Ano93a]. California [IEE93]. Canada [BH95]. Carmel [CHK^{+}93a, CHK^{+}93b]. Case [PS96]. Chambery [Ral94]. Chapter [Saf98c]. Clauses [SU92, SU94a, SU94b]. Comparing [SPU94]. Comparison [CHK^{+}93a, CHK^{+}93b]. Conference [Ano93a, Ano93b, BH95, CKM93, IEE93, LS85, VY93]. Construction [SU92, SU94a, SU94b]. Control [SRK97, Saf99, SRK93b, SKR95, Saf98c, Saf98b]. Controller [RKS93, SRK93a, SHK^{+}93]. coordination [Saf97a].

Data [CS99]. DC [Ano93b]. Debugging [PS96]. December [DvLV96]. Default [BSS95a, BSS95b, BSS00]. Dempster [Saf91b, Saf94a, YKF94]. Dempster-Shafer [Saf91b, YKF94]. Dempster-Shafer’s [Saf94a]. Design [KMRS97]. determination [SS97]. Detroit [IEE95]. Dialogue [SS88b]. Directedness [SRK93a]. distributed [PS93]. Domain [JS02]. Doubtful [Saf91a]. Driven [SU94a, SU94b].

ECSQARU [CKM93]. education [JPD96]. Electric [JS02]. Eleventh [Ano93b, BH95]. European [CKM93]. evidence [YKF94]. execution [SRK94].

Field [JS02]. First [SU94a, SU94b]. First-Order [SU94a, SU94b]. Flakey [CHK^{+}93a, CHK^{+}93b, RKS93, SHK^{+}93]. formalisms [PS93]. Framework [Saf90, BDL^{+}98]. France [JPD96, Ral94]. Francisco [IEE93]. Function [Saf92]. Functions [BSS95a, BSS95b, BSS00]. fusing [BS99]. fusion [SS97]. Fuzzy [Ano93a, CS99, IEE93, RKS93, SRK93a, SHK^{+}93, Saf97a, Saf97b, Saf97c, SRK97, Saf98a, SRK93b, SRK93c, SRK94, SW95a, SW96, SS97, YLZ95, Ral94]. fuzzy-based [SW95a].

General [SU91, SU92]. Goal [SRK93a]. Goal-Directedness [SRK93a]. Granada [CKM93].

Handling [Saf98c, Saf98b, Saf99, PS93]. Hierarchical [SW95a, SW95b]. Human [CKS10]. Human-aware [CKS10]. Hybrid [SS88b, Saf90, Saf91a, SS89].

IJCAI [Ral94]. implicitly [BDL^{+}98].

implicitly-specified [BDL^{+}98].

inconsistency [BDL^{+}98]. Industrial [CCM^{+}85, YLZ95]. Inference [SU94a, SU94b]. Inference-Driven [SU94a, SU94b]. Information [PS95a, VY93]. Integrating [PS93, SKR95]. Integration [Saf93]. Intelligence [Ano93b, BHKL91, LS85, BH95, PS93, Ral94, VY93]. Intelligent [IEE95, YLZ95]. Intentions [PPSW99]. International [IEE93, JPD96, LS85, VY93, DvLV96]. IPMU [VY93]. ISRAM [JPD96]. Issues [Saf94a]. Italy [LS85].

July [Ano93a, Ano93b, VY93].

Knowledge [Saf90, Saf91b, VY93, Saf94a]. Knowledge-Based [VY93]. Krypton [SS88b].

Language [SS88b]. localization

[SW95a, SW96, Saf96]. locations [SW96]. locative [SW95b]. Logic [Ano93a, CS99, Saf92, Saf97b, Saf97c, SRK97, Saf98a, BS99, Ral94, SS89, SRK93b, SRK93c, SRK94, SKR95, Saf97a, YLZ95].

M [SS88b]. M-Krypton [SS88b]. Mallorca [VY93]. Management [SPU94, VY93, PS95a]. manufacturing [JPD96]. March [IEE93]. May [JPD96]. McGill [BH95]. Medicine [LS85]. methods [VY93]. Mobile [RKS93, Saf93, SRK97, CKS10, SRK93b, SW95a, SW95b]. modal [BS99]. Model [SHS^{+}91]. Modelling [SS88b]. Monoclonal [CCM^{+}85]. Montpellier [JPD96]. Montreal [BH95]. motion [RSK95, SRK93c]. Multiple [SS88b, BS99]. multivalued [SKR95]. multivalued-logic [SKR95].

National [Ano93b]. Navigation [Saf98a, Saf96]. Netherlands [DvLV96]. Networks [SU91]. Notes [Saf93]. November [CKM93]. Numerical [PS96].

online [Saf97b]. Order [SU94a, SU94b].

Palma [VY93]. partial [BDL^{+}98, BS99]. Pavia [LS85]. Perception [SW96]. Perception-based [SW96]. perceptual [KBB^{+}08]. Pick [Saf94b]. Pick-up [Saf94b]. Pitch [SS97]. Planning [Saf93, CKS10, KBB^{+}08, RSK95, SRK93c, SKR95]. plans [SRK94]. platforms [SW95b]. point [Saf88]. probability [BDL^{+}98]. Proceedings [Ano93b, IEE95, BH95, CKM93, DvLV96, JPD96, LS85, VY93, Ral94]. Processing [VY93, PS95a]. Production [CCM^{+}85]. Progress [RSK95]. Propagating [SU91]. PULCINELLA [SU91].

Qualitative [PS95b, PS96, BDL^{+}98]. Quantitative [PS95b, CKM93]. Quebec [BH95].

Reactivity [SRK93a, Saf93]. reasoners [BS99]. Reasoning [BSS95a, BDL^{+}98, DvLV96, BSS95b, BSS00, CKM93, SW95b]. recovery [KBB^{+}08]. relations [BDL^{+}98]. Representation [Saf91b, SS88a, Saf94a]. Representing [Saf90]. research [JPD96, RSK95]. RoboCup [JS02]. Robot [RKS93, SHK^{+}93, Saf96, SRK97, Saf98a, SRK93b, SRK94]. Robotics [JPD96, Saf97b, Saf97c, DvLV96, Saf97a]. Robots [PPSW99, Saf93, Saf99, CKS10, SW95a, Saf98c, Saf98b]. Robust [SRK93b, SRK94]. RUR [DvLV96].

San [IEE93]. Saphira [KMRS97]. Second [IEE93]. secure [KBB^{+}08]. self [SW96, Saf96]. self-localization [SW96, Saf96]. September [IEE95, LS85]. Shafer [Saf91b, YKF94]. Shafer’s [Saf94a]. signals [SS97]. Sixth [JPD96]. Skilled [CCM^{+}85]. Some [Saf93]. Spain [CKM93, VY93]. specified [BDL^{+}98]. speech [SS97]. Study [PS96]. Sweden [SBB^{+}01, SBJW02]. Symbolic [CKM93]. Symbols [CS99]. Symposium [IEE95, JPD96]. System [Saf91a]. Systems [IEE93, SU92, SU94a, SU94b, VY93, YLZ95].

task [CKS10]. Team [SBB^{+}01, SBJW02]. Techniques [SPU94]. Theory [Saf91b, Saf94a, YKF94]. there [Saf88]. Tool [SU91]. Transferable [SHS^{+}91]. Treatment [Saf87, Saf88]. trends [JPD96]. Two [CHK^{+}93a, CHK^{+}93b].

Uncertain [Saf90]. Uncertainty

[BH95, BHKL91, PS95b, PS96, Saf87, SU91, SPU94, Saf99, VY93, CKM93, DvLV96, PS93, PS95a, Saf88, Saf98c, Saf98b]. unified [BDL^{+}98]. University [BH95]. USA [IEE95]. Using [JS02, Saf91b, SRK93c, Saf97b, Saf97c, SRK97, SRK93b, SRK94, SW96].

Valuation [SU91, SU92, SU94a, SU94b]. vantage [Saf88]. vehicle [RSK95, SRK93c]. Vehicles [IEE95]. Verification [PS95b, PS96]. View [Saf87]. Vision [CS99]. vs [CHK^{+}93a, CHK^{+}93b].

Washington

[Ano93b]. Winners [CHK^{+}93a, CHK^{+}93b]. workshop [DvLV96, Ral94, Saf97b].

References

Anonymous:1993:FLC

[Ano93a] Anonymous, editor. Fuzzy Logic ’93: Conference — July 1993, Burlingame, CA. Computer Design, Westford, MA, USA, 1993.

Anonymous:1993:PEN

[Ano93b] Anonymous, editor. Proceedings of the Eleventh National Conference on Artificial Intelligence, July 11–15, 1993, Washington, DC. AAAI Press and MIT Press, Menlo Park, CA, USA and Cambridge, MA, USA, 1993. ISBN 0-262-51071-5. LCCN Q 334 N37 1993.

Benferhat:1998:RUI

[BDL^{+}98] S. Benferhat, D. Dubois, J. Lang, H. Prade, A. Saffiotti, and P. Smets. Reasoning under inconsistency based on implicitlyspecified partial qualitative probability relations: a unified framework. In AAAI-98/IAAI-98 Proceedings (Madison, WI, 1998), pages 121–126. MIT Press, Cambridge, MA, 1998.

Besnard:1995:UAI

[BH95] Philippe Besnard and Steve Hanks, editors. Uncertainty in artificial intelligence: proceedings of the Eleventh Conference (1995): August 18–20, 1995 / Eleventh Conference on Uncertainty in Artificial Intelligence, McGill University, Montreal, Quebec, Canada. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1995. ISBN 1-55860-385-9. LCCN Q375 .C66 1995.

Bonissone:1991:ACV

[BHKL91] Piero Patrone Bonissone, M. Henrion, L. N. Kanal, and J. Lemmer, editors. Uncertainty in Artificial Intelligence 6, volume 12 of Machine intelligence and pattern recognition. Elsevier Science Publishers, Amsterdam, The Netherlands, 1991. ISBN 0-444-89264-8. LCCN Q335.U536 1991. Papers presented at the Sixth Conference on Uncertainty in Artificial Intelligence, held in July, 1990, at the Massachusetts Institute of Technology, Cambridge, MA.

Boldrin:1999:MLF

[BS99] Luca Boldrin and Alessandro Saffiotti. A modal logic for fusing partial belief of multiple reasoners. J. Logic Comput., 9(1):81– 103, 1999. ISSN 0955-792X.

Benferaht:1995:BFD

[BSS95a] S. Benferaht, A. Saffiotti, and P. Smets. Belief functions and default reasoning. In Besnard and Hanks [BH95], pages 19–26. ISBN 1-55860-385-9. LCCN Q375 .C66 1995.

                                        URL ftp://iridia.ulb.ac.be/
                                        pub/saffiotti/robot/uai95.
                                        ps;   http://iridia.ulb.ac.
                                        be/saffiotti/abstracts.html#

uai95. Extended version available as TR/IRIDIA/9505.

Benferhat:1995:BFD

[BSS95b] S. Benferhat, A. Saffiotti, and P. Smets. Belief functions and default reasoning. In Uncertainty in artificial intelligence (Montréal, PQ, 1995), pages 19–26. Morgan Kaufmann, San Francisco, CA, 1995.

Benferhat:2000:BFD

[BSS00] S. Benferhat, A. Saffiotti, and P. Smets. Belief functions and default reasoning. Artificial Intelligence, 122(1-2):1–69, 2000. CO-DEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Cerri:1985:TSA

[CCM^{+}85] S. A. Cerri, A. Cherubini, C. Minciotti, A. Saffiotti, and R. Sbarbati. Towards a skilled assistant in the industrial production of monoclonal antibodies. In Lotto and Stefanelli [LS85], pages 95– 105. ISBN 0-444-87843-2. LCCN R859.7.D42 I57 1985.

Congdon:1993:CVF

[CHK^{+}93a] C. Congdon, M. Huber, D. Kortenkamp, K. Konolige, K. Myers, E. H. Ruspini, and A. Saffiotti. Carmel vs. Flakey: A comparison of two winners. The AI Magazine, 14(1):49–57, Spring 1993. CODEN AIMAEK. ISSN 0738-4602. URL http://

iridia.ulb.ac.be/saffiotti/
abstracts.html#aim93.

Congdon:1993:SCV

[CHK^{+}93b] C. Congdon, M. Huber, D. Kortenkamp, K. Konolige, K. Myers, E. H. Ruspini, and A. Saffiotti. Carmel vs. Flakey: A comparison of two winners. The AI Magazine, 14(1):49–57, Spring 1993. CODEN AIMAEK. ISSN 0738-4602.

Clarke:1993:SQA

[CKM93] Michael Clarke, Rudolf Kruse, and Serafin Moral, editors. Symbolic and quantitative approaches to reasoning and uncertainty: European Conference ECSQARU ’93, Granada, Spain, November 8–10, 1993: proceedings, volume 747 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNCSD9. ISBN 0-387-57395-X (USA), 3-540-57395-X (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q375 .E88 1993. DM72.00.

Cirillo:2010:HAT

[CKS10] Marcello Cirillo, Lars Karlsson, and Alessandro Saffiotti. Humanaware task planning: An application to mobile robots. ACM Transactions on Intelligent Systems and Technology (TIST), 1 (2):15:1–15:??, November 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Coradeschi:1999:ASV

[CS99] Silvia Coradeschi and Alessandro Saffiotti. Anchoring symbols to vision data by fuzzy logic. Lecture Notes in Computer Science, 1638:104–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/1638/16380104.htm;
 http://link.springer-ny.com/
link/service/series/0558/papers/
1638/16380104.pdf.

Dorst:1996:RUR

[DvLV96] Leo Dorst, Michiel van Lambalgen, and Frans Voorbraak, editors. Reasoning with uncertainty in robotics: international workshop, RUR ’95, Amsterdam, The Netherlands, December 4–6, 1995: proceedings, volume 1093 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61376-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ210.3 .R87 1995. Contents: Mathematical foundations of navigation and perception for an autonomous mobile robot / J.L. Crowley – Reasoning with uncertainty in AI / F. Voorbraak – Robot navigation: integrating perception, environmental constraints and task execution within a probabilistic framework / A. Elfes – Uncertainty reasoning in object recognition by image processing / R. Hummel – Partially observable Markov decision processes for artificial intelligence / L.P. Kaelbling, M.L. Littman, A.R. Cassandra – An evidential approach to probabilistic mapbuilding / D. Pagac, E.M. Nebot, H. Durrant-Whyte – Belief formation by constructing models / J. Paris, A. Vencovska – Causal relevance / G. Shafer – The robot control strategy in a domain with dynamical obstacles / G.V. Alferov, O.A. Malafeyev – Reasoning about noisy sensors (and effectors) in the situation calculus / F. Bacchus, J.Y. Halpern, H.J. Levesque – Recursive total least squares: an alternative to using the discrete Kalman filter in robot navigation / D.L. Boley, E.S. Steinmetz, K.T. Sutherland – A sensor-based motion planner for mobile robot navigation with uncertainty / B. Bouilly, T. Simeon – Knowledge considerations in robotics / R.I. Brafman, Y. Shoham – Neural network applications in sensor fusion for an autonomous mobile robot / J.W.M. van Dam, B.J.A. Krose, F.C.A. Groen – Structuring uncertain knowledge with hierarchical Bayesian networks / T. Frohlinghaus – Uncertainty treatment in a surface filling mobile robot / E. Gonzalez ... [et al.] – Probabilistic map learning: necessity and difficulties / P. Hebert, S. Betge-Brezetz, R. Chatila – Robot navigation with Markov models: a framework for path planning and learning with limited computational resources / S. Koenig, R. Goodwin, R.G. Simmons – A refined method for occupancy grid interpretation / K. Konolige – Sensor planning with Bayesian decision theory / S. Kristensen – Perception-based self-localization using fuzzy locations / A. Saffiotti, L.P. Wesley.

IEEE:1993:SII

[IEE93] IEEE, editor. Second IEEE International Conference on Fuzzy Systems: San Francisco, California, March 28–April 1, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-0614-7 (paperback), 0-7803-0615-5 (case), 0-7803-0616-3 (microfiche), 0-7803-0615-5 (casebound), 0-7803-0614-7 (softbound). LCCN TJ212.2.I3249 1993. Two volumes. IEEE catalog number: 93CH3136-9.

IEEE:1995:PIV

[IEE95] IEEE, editor. Proceedings of the Intelligent Vehicles ’95 Symposium, September 25–26, 1995, Detroit, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2983-X (softbound), 0-7803-2984-8 (microfiche). LCCN TL272.5 .I62 1995. IEEE catalog number 95TH8132.

Jamshidi:1996:RMR

[JPD96] Mohammad Jamshidi, François Pin, and Pierre Dauchez, editors. Robotics and manufacturing: recent trends in research, education, and applications: proceedings of the Sixth International Symposium on Robotics and Manufacturing (ISRAM ’96), May 28–30, 1996, Montpellier France, volume 6 of ASME Press series—proceedings of the international symposia on robotics and manufacturing. ASME Press, New York, NY, USA, 1996. ISBN 0-7918-0047-4. ISSN 1052-4150. LCCN TS191.8.I58 1996.

Johansson:2002:UEF

[JS02] Stefan J. Johansson and Alessandro Saffiotti. Using the electric field approach in the RoboCup domain. Lecture Notes in Computer Science, 2377:399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-
                                        ny.com/link/service/series/
                                        0558/bibs/2377/23770399.htm;
                                         http://link.springer-ny.com/
                                        link/service/series/0558/papers/
                                        2377/23770399.pdf.

Karlsson:2008:SAR

[KBB^{+}08] Lars Karlsson, Abdelbaki Bouguerra, Mathias Broxvall, Silvia Coradeschi, and Alessandro Saffiotti. To secure an anchor—a recovery planning approach to ambiguity in perceptual anchoring. AI Commun., 21(1):1–14, 2008. ISSN 0921-7126.

Konolige:1997:SAD

[KMRS97] K. Konolige, K. L. Myers, E. H. Ruspini, and A. Saffiotti. The Saphira architecture: A design for autonomy. Journal of experimental & theoretical artificial intelligence: JETAI, 9 (1):215–235, 1997. CODEN JEAIEL. ISSN 0952-813X. URL

                                        ftp://iridia.ulb.ac.be/pub/
                                        saffiotti/robot/jetai96.ps;
                                         http://iridia.ulb.ac.be/saffiotti/
                                        abstracts.html#jetai96.

Lotto:1985:AIM

[LS85] Ivo De Lotto and Mario Stefanelli, editors. Artificial Intelligence in Medicine: proceedings of the International Conference on Artificial Intelligence in Medicine, Pavia, Italy, 13– 14 September 1985. North-Hol-land, Amsterdam, The Netherlands, 1985. ISBN 0-444-87843-2. LCCN R859.7.D42 I57 1985.

Parsons:1999:RBI

[PPSW99] Simon Parsons, Ola Pettersson, Alessandro Saffiotti, and Michael Wooldridge. Robots with the best of intentions. Lecture Notes in Computer Science, 1600:329–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/1600/16000329.htm;
http://link.springer-ny.com/
link/service/series/0558/papers/
1600/16000329.pdf.

Parsons:1993:IUH

[PS93] S. Parsons and A. Saffiotti. Integrating uncertainty handling formalisms in distributed artificial intelligence. In Clarke et al. [CKM93], pages 304–309. CO-DEN LNCSD9. ISBN 0-387-57395-X (USA), 3-540-57395-X (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q375 .E88 1993. DM72.00.

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/uncertainty/
lncs747.ps;   http://iridia.
ulb.ac.be/saffiotti/abstracts.
html#lncs747.

Parsons:1995:IPM

[PS95a] S. Parsons and A. Saffiotti. Information processing and the management of uncertainty. The Knowledge Engineering Review, 10(1):83–88, 1995. CODEN KEREE3. ISSN 0269-8889. Annotated conference report.

Parsons:1995:QVQ

[PS95b] S. Parsons and A. Saffiotti. The qualitative verification of quantitative uncertainty. Lecture Notes in Computer Science, 945:180– 189, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). A preliminary version appeared in the Procs. of the IPMU Conference, Paris, France, 1994.

Parsons:1996:CSQ

[PS96] S. Parsons and A. Saffiotti. A case study in the qualitative verification and debugging of numerical uncertainty. International Journal of Approximate Reasoning, 14(2/3):187–216, 1996. CO-DEN IJARE4. ISSN 0888-613X.

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/uncertainty/
ijar96.ps;   http://iridia.
ulb.ac.be/saffiotti/abstracts.
html#ijar96.

Ralescu:1994:FLA

[Ral94] Anca L. Ralescu, editor. Fuzzy logic in artificial intelligence: IJ-CAI ’93 workshop, Chambery, France, August 28, 1993: proceedings, volume 847 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 3-540-58409-9 (Berlin),, 0-387-58409-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QC173.454 .F89 1994. DM39.00.

Ruspini:1993:FCF

[RKS93] E. Ruspini, K. Konolige, and A. Saffiotti. A fuzzy controller for Flakey, an autonomous mobile robot. In Anonymous [Ano93a], page A213.

Ruspini:1995:PRA

[RSK95] E. H. Ruspini, A. Saffiotti, and K. Konolige. Progress in research on autonomous vehicle motion planning. In Yen et al. [YLZ95], pages xviii + 356. ISBN 0-7803-1048-9. LCCN TJ213 .I5376 1995. URL

ftp://iridia.ulb.ac.be/pub/
saffiotti/robot/yen_book95.
ps;   http://iridia.ulb.ac.
be/saffiotti/abstracts.html#

yen_book. Chapter 8.

Saffiotti:1987:AVT

[Saf87] A. Saffiotti. An AI view of the treatment of uncertainty. The Knowledge Engineering Review, 2(2):75–97, 1987. CO-DEN KEREE3. ISSN 0269-

8889. URL http://iridia.ulb.
ac.be/saffiotti/abstracts.

html#ker87. See the panel discussion about this paper on KER 3(1) 59–86.

Saffiotti:1988:TUA

[Saf88] A. Saffiotti. The treatment of uncertainty in AI: is there a better vantage point? The Knowledge Engineering Review, 3(1):87–91, 1988. CODEN KEREE3. ISSN 0269-8889. Rebuttal in panel discussion.

Saffiotti:1990:HFR

[Saf90] A. Saffiotti. A hybrid framework for representing uncertain knowledge. In Procs. of the AAAI Conf., pages 653–658. AAAI Press, Menlo Park, California, 1990. URL

ftp://iridia.ulb.ac.be/pub/
saffiotti/uncertainty/aaai90.
ps;   http://iridia.ulb.ac.
be/saffiotti/abstracts.html#
aaai90.

Saffiotti:1991:HBS

[Saf91a] A. Saffiotti. A hybrid belief system for doubtful agents. Lecture Notes in Computer Science, 521:393–402, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/uncertainty/
lncs251.ps;   http://iridia.
ulb.ac.be/saffiotti/abstracts.
html#lncs251.

Saffiotti:1991:UDS

[Saf91b] A. Saffiotti. Using Dempster-Shafer theory in knowledge representation. In Bonissone et al. [BHKL91], pages 417–431. ISBN 0-444-89264-8. LCCN Q335.U536 1991. Papers presented at the Sixth Conference on Uncertainty in Artificial Intelligence, held in July, 1990, at the Massachusetts Institute of Technology, Cambridge, MA.

Saffiotti:1992:BFL

[Saf92]  Alessandro Saffiotti.  A be-

lief function logic. In AAAI-92. Proceedings, Tenth National Conference on Artificial Intelligence (San Jose, CA, 1992), pages 642–647. AAAI Press, Menlo Park, CA, USA, 1992.

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/uncertainty/
aaai92.ps;   http://iridia.
ulb.ac.be/saffiotti/abstracts.
html#aaai92.

Saffiotti:1993:SNI

[Saf93] A. Saffiotti. Some notes on the integration of planning and reactivity in autonomous mobile robots. In Procs. of the AAAI Spring Symposium on Foundations of Automatic Planning, pages 122–126. AAAI Press, Menlo Park, California,

1993. URL ftp://iridia.ulb.
ac.be/pub/saffiotti/robot/
aaai_sss93.ps;   http://iridia.
ulb.ac.be/saffiotti/abstracts.
html#sss93.

Saffiotti:1994:IKR

[Saf94a] A. Saffiotti. Issues of knowledge representation in Dempster-Shafer’s theory. In Yager et al. [YKF94], pages 415– 440. ISBN 0-471-55248-8. LCCN QA76.87 .A39 1994.

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/uncertainty/
yager_book94.ps;   http://
iridia.ulb.ac.be/saffiotti/
abstracts.html#yager_book.

Saffiotti:1994:PW

[Saf94b] A. Saffiotti. Pick-up what? In C. Bäckström and E. Sandewall, editors, Current Trends in AI Planning — Procs. of EWSP ’93, pages 266–277. IOS Press, Amsterdam, The Netherlands, 1994.

                                        URL ftp://iridia.ulb.ac.be/
                                        pub/saffiotti/robot/ewsp93.
                                        ps;   http://iridia.ulb.ac.
                                        be/saffiotti/abstracts.html#
                                        ewsp93.

Saffiotti:1996:RNU

[Saf96]  A. Saffiotti.  Robot naviga-

tion under approximate selflocalization. In Jamshidi et al. [JPD96], pages 589–594. ISBN 0-7918-0047-4. ISSN 1052-4150. LCCN TS191.8.I58 1996.

                                        URL ftp://iridia.ulb.ac.be/
                                        pub/saffiotti/robot/wac96.
                                        ps;   http://iridia.ulb.ac.
                                        be/saffiotti/abstracts.html#
                                        wac96.

Saffiotti:1997:FLA

[Saf97a] A. Saffiotti. Fuzzy logic in autonomous robotics: behavior coordination. In Procs. of the 6th IEEE Int. Conf. on Fuzzy Systems, pages 573–578. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

                                        1997. URL ftp://iridia.ulb.
                                        ac.be/pub/saffiotti/robot/
                                        ieee_fuzzy97.ps;   http://
                                        iridia.ulb.ac.be/saffiotti/
                                        abstracts.html#ieee97.

Saffiotti:1997:UFLa

[Saf97b] A. Saffiotti. Using fuzzy logic for autonomous robotics — an online workshop. The Knowledge Engineering Review, 12 (1):91–94, 1997. CODEN KEREE3. ISSN 0269-8889.

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/robot/ker97.
ps;   http://iridia.ulb.ac.
be/saffiotti/abstracts.html#

ker97. Annotated conference report.

Saffiotti:1997:UFLb

[Saf97c] A. Saffiotti. Using fuzzy logic in autonomous robotics. Soft computing: newsletter of the Canadian Society for Fuzzy Information and Neural Systems, 1 (4):??, 1997. ISSN 1195-3934.

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/robot/tr9706.
ps;   http://iridia.ulb.ac.
be/FLAR/survey.html. To ap-

pear.

Saffiotti:1998:FLA

[Saf98a] A. Saffiotti. Fuzzy logic in autonomous robot navigation. In P. Bonissone E. H. Ruspini and W. Pedrycz, editors, IEEE Handbook of Fuzzy Computation. Oxford University Press, Oxford, UK, 1998. Forthcoming.

Saffiotti:1998:HUC

[Saf98b] A. Saffiotti. Handling uncertainty in control of autonomous robots. In A. Hunter and S. Parsons, editors, Uncertainty in Information Systems, Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

1998. URL ftp://iridia.ulb.
ac.be/pub/saffiotti/robot/

tr9708.ps. Forthcoming.

Saffiotti:1998:CHU

[Saf98c] Alessandro Saffiotti. Chapter 10: Handling uncertainty in control of autonomous robots. Lecture Notes in Computer Science, 1455:198–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-
                                        ny.com/link/service/series/
                                        0558/bibs/1455/14550198.htm;
                                         http://link.springer-ny.com/
                                        link/service/series/0558/papers/
                                        1455/14550198.pdf.

Saffiotti:1999:HUC

[Saf99]  Alessandro Saffiotti.  Han-

dling uncertainty in control of autonomous robots. Lecture Notes in Computer Science, 1600:381–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-
                                        ny.com/link/service/series/
                                        0558/bibs/1600/16000381.htm;
                                         http://link.springer-ny.com/
                                        link/service/series/0558/papers/
                                        1600/16000381.pdf.

Saffiotti:2001:TS

[SBB^{+}01] A. Saffiotti, M. Boman, P. Buschka, P. Davidsson, S. Johansson, and Z. Wasik. Team sweden. Lecture Notes in Computer Science, 2019:643–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-
                                        ny.com/link/service/series/
                                        0558/bibs/2019/20190643.htm;
                                         http://link.springer-ny.com/

link/service/series/0558/papers/
2019/20190643.pdf.

Saffiotti:2002:TS

[SBJW02] A. Saffiotti, A. Björklund, S. Johansson, and Z. Wasik. Team sweden. Lecture Notes in Computer Science, 2377:725– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer-ny.com/link/
service/series/0558/bibs/2377/
23770725.htm;   http://link.
springer-ny.com/link/service/
series/0558/papers/2377/23770725.
pdf.

Saffiotti:1993:FCF

[SHK^{+}93] A. Saffiotti, N. Helft, K. Konolige, J. Lowrance, K. Myers, D. Musto, E. Ruspini, and L. Wesley. A fuzzy controller for Flakey, the robot. In Anonymous [Ano93b], page 864. ISBN 0-262-51071-5. LCCN Q 334 N37 1993. Video abstract.

Smets:1991:TBM

[SHS^{+}91] P. Smets, Y-T Hsia, A. Saffiotti, R. Kennes, H. Xu, and E. Umkehrer. The transferable belief model. Lecture Notes in Computer Science, 548:91–98, October 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saffiotti:1995:MLA

[SKR95] A. Saffiotti, K. Konolige, and E. H. Ruspini. A multivaluedlogic approach to integrating planning and control. Artificial Intelligence, 76(1-2):481– 526, 1995. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic). URL ftp://

iridia.ulb.ac.be/pub/saffiotti/
robot/aij95.ps;   http://iridia.
ulb.ac.be/saffiotti/abstracts.
html#aij95.

Saffiotti:1994:CUM

[SPU94] A. Saffiotti, S. Parsons, and E. Umkehrer. Comparing uncertainty management techniques. Microcomputers in Civil Engineering, 9:367–380, 1994. CO-DEN MCENE7. ISSN 0885-

9507. URL http://iridia.ulb.
ac.be/saffiotti/abstracts.
html#mce94.

Saffiotti:1993:BRG

[SRK93a] A. Saffiotti, E. H. Ruspini, and K. Konolige. Blending reactivity and goal-directedness in a fuzzy controller. In IEEE [IEE93], pages 134–139. ISBN 0-7803-0614-7 (paperback), 0-7803-0615-5 (case), 0-7803-0616-3 (microfiche), 0-7803-0615-5 (casebound), 0-7803-0614-7 (softbound). LCCN TJ212.2.I3249

1993. URL ftp://iridia.ulb.
ac.be/pub/saffiotti/robot/
ieee_fuzzy93.ps;   http://
iridia.ulb.ac.be/saffiotti/
abstracts.html#ieee93. Two

volumes. IEEE catalog number: 93CH3136-9.

Saffiotti:1993:RCM

[SRK93b] A. Saffiotti, E. H. Ruspini, and K. Konolige. Robust control of a mobile robot using fuzzy logic. In Procs. of the First European Congress on Fuzzy and Intelligent Technologies (EUFIT), pages 881–886. ELITE Foundation, Aachen, Germany, 1993.

Saffiotti:1993:UFL

[SRK93c] A. Saffiotti, E. H. Ruspini, and K. Konolige. Using fuzzy logic for autonomous vehicle motion planning. In Proceedings of the NAFIPS Conference. ????, Allentown, PA, 1993.

Saffiotti:1994:RER

[SRK94] A. Saffiotti, E. H. Ruspini, and K. Konolige. Robust execution of robot plans using fuzzy logic. In Ralescu [Ral94], pages 24–

37. CODEN LNCSD9. ISBN 3-540-58409-9 (Berlin),, 0-387-58409-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QC173.454 .F89 1994. DM39.00. URL

ftp://iridia.ulb.ac.be/pub/
saffiotti/robot/lnai847.ps;
 http://iridia.ulb.ac.be/saffiotti/
abstracts.html#lnai847.

Saffiotti:1997:UFLc

[SRK97] A. Saffiotti, E. H. Ruspini, and K. Konolige. Using fuzzy logic for mobile robot control. In H. Prade D. Dubois and H. J. Zimmermann, editors, International Handbook of Fuzzy Sets and Possibility Theory, volume 5. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997.

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/robot/hfs97.

                                        ps;   http://iridia.ulb.ac.
                                        be/saffiotti/abstracts.html#

hfs97. Forthcoming.

Saffiotti:1988:AAB

[SS88a] A. Saffiotti and F. Sebastiani. Acquaintance and awareness in belief representation. In Procs. of the Computational Intelligence Conf., pages 159–161. Elsevier Science Publishers, Amsterdam, The Netherlands, 1988.

Saffiotti:1988:DMM

[SS88b] A. Saffiotti and F. Sebastiani. Dialogue modelling in M-krypton, a hybrid language for multiple believers. In Proceedings of the 4th IEEE Conference on Artificial Intelligence Applications, pages 56–60. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. URL http://

                                        iridia.ulb.ac.be/saffiotti/
                                        abstracts.html#caia88.

Saffiotti:1989:THL

[SS89] A. Saffiotti and F. Sebastiani. Towards a hybrid logic of acquaintance and awareness. In A. Cohn, editor, Procs. of the AISB 89 Conf. on AI, pages 95– 103. Pitman Publishing, London, UK, 1989. URL http://

                                        iridia.ulb.ac.be/saffiotti/
                                        abstracts.html#aisb89.

Saffiotti:1997:PDS

[SS97] A. Saffiotti and A. Soquet. Pitch determination of speech signals: a fuzzy fusion approach. In Procs. of the 7th IFSA World Congress, volume I, pages 315–320. Academia Press, Prague, CK, 1997. URL

ftp://iridia.ulb.ac.be/pub/
saffiotti/robot/ifsa97.ps;
 http://iridia.ulb.ac.be/saffiotti/
abstracts.html#ifsa97.

Saffiotti:1991:PGT

[SU91] A. Saffiotti and E. Umkehrer.

PULCINELLA: A general tool for propagating uncertainty in valuation networks. In Procs. of the 7th Conf. on Uncertainty in AI, pages 323–331. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1991.

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/uncertainty/
uai91.ps;   http://iridia.
ulb.ac.be/saffiotti/abstracts.
html#uai91.

Saffiotti:1992:ACV

[SU92] A. Saffiotti and E. Umkehrer.

Automatic construction of valuation systems from general clauses. In Valverde and Yager [VY93], pages 341–345. CODEN LNCSD9. ISBN 0-387-56735-6 (New York), 3-540-56735-6 (Berlin: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I557 1992. US$54.00, DM72.00.

Saffiotti:1994:ACV

[SU94a] A. Saffiotti and E. Umkehrer.

Inference-driven construction of valuation systems from first-order clauses. IEEE Trans. Systems, Man, and Cybernetics, 24(11): 1611–1624, 1994. CODEN ISY-MAW. ISSN 0018-9472, 1083-4427.

Saffiotti:1994:IDC

[SU94b] A. Saffiotti and E. Umkehrer. Inference-driven construction of valuation systems from firstorder clauses. IEEE Trans. Systems, Man, and Cybernetics, 24(11):1611–1624, 1994. CO-DEN ISYMAW. ISSN 0018-9472, 1083-4427. URL http://

                                        iridia.ulb.ac.be/saffiotti/
                                        abstracts.html#smc94.

Saffiotti:1995:HFB

[SW95a] A. Saffiotti and L. P. Wesley. Hierarchical fuzzy-based localization in autonomous mobile robots. In Procs. of the IFSA Conference. ????, San Paulo, Brazil, 1995.

Saffiotti:1995:HLR

[SW95b] A. Saffiotti and L. P. Wesley. Hierarchical locative reasoning for autonomous mobile platforms. In IEEE [IEE95], pages 202–207. ISBN 0-7803-2983-X (softbound), 0-7803-2984-8 (microfiche). LCCN TL272.5 .I62

                                        1995. URL http://iridia.ulb.
                                        ac.be/saffiotti/abstracts.

html#iv95. IEEE catalog number 95TH8132.

Saffiotti:1996:PBS

[SW96] A. Saffiotti and L. P. Wesley. Perception-based selflocalization using fuzzy locations. In Dorst et al. [DvLV96], pages 368–385. ISBN 3-540-61376-5. ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN TJ210.3 .R87 1995.

URL ftp://iridia.ulb.ac.be/
pub/saffiotti/robot/rur96.
ps;   http://iridia.ulb.ac.
be/saffiotti/abstracts.html#

rur96. Contents: Mathematical foundations of navigation and perception for an autonomous mobile robot / J.L. Crowley – Reasoning with uncertainty in AI / F. Voorbraak – Robot navigation: integrating perception, environmental constraints and task execution within a probabilistic framework / A. Elfes – Uncertainty reasoning in object recognition by image processing / R. Hummel – Partially observable Markov decision processes for artificial intelligence / L.P. Kaelbling, M.L. Littman, A.R. Cassandra – An evidential approach to probabilistic mapbuilding / D. Pagac, E.M. Nebot, H. Durrant-Whyte – Belief formation by constructing models / J. Paris, A. Vencovska – Causal relevance / G. Shafer – The robot control strategy in a domain with dynamical obstacles / G.V. Alferov, O.A. Malafeyev – Reasoning about noisy sensors (and effectors) in the situation calculus / F. Bacchus, J.Y. Halpern, H.J. Levesque – Recursive total least squares: an alternative to using the discrete Kalman filter in robot navigation / D.L. Boley, E.S. Steinmetz, K.T. Sutherland – A sensor-based motion planner for mobile robot navigation with uncertainty / B. Bouilly, T. Simeon – Knowledge considerations in robotics / R.I. Brafman, Y. Shoham – Neural network applications in sensor fusion for an autonomous mobile robot / J.W.M. van Dam, B.J.A. Krose, F.C.A. Groen – Structuring uncertain knowledge with hierarchical Bayesian networks / T. Frohlinghaus – Uncertainty treatment in a surface filling mobile robot / E. Gonzalez ... [et al.] – Probabilistic map learning: necessity and difficulties / P. Hebert, S. Betge-Brezetz, R. Chatila – Robot navigation with Markov models: a framework for path planning and learning with limited computational resources / S. Koenig, R. Goodwin, R.G. Simmons – A refined method for occupancy grid interpretation / K. Konolige – Sensor planning with Bayesian decision theory / S. Kristensen – Perception-based self-localization using fuzzy locations / A. Saffiotti, L.P. Wesley.

Valverde:1993:IAM

[VY93] Llorenc Valverde and Ronald R. Yager, editors. IPMU ’92, advanced methods in artificial intelligence: 4th International Conference on Information Processing and Management of Uncertainty in Knowledge-Based Systems, Palma de Mallorca, Spain, July 6–10, 1992: proceedings, volume 682 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNCSD9. ISBN 0-387-56735-6 (New York), 3-540-56735-6 (Berlin: paperback).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I557 1992. US$54.00, DM72.00.

Yager:1994:ADS

[YKF94] Ronald R. Yager, Janusz Kacprzyk, and Mario Fedrizzi, editors. Advances in the Dempster-Shafer theory of evidence. Wiley, New York, NY, USA, 1994. ISBN 0-471-55248-8. vii + 597 pp. LCCN QA76.87 .A39 1994.

Yen:1995:IAF

[YLZ95] John Yen, Reza Langari, and Lotfi A. Zadeh, editors. Industrial applications of fuzzy logic and intelligent systems. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-1048-9. LCCN TJ213 .I5376